### v1

# On the Invariant and Geometric Structure of the Holomorphic Unified Field Theory

Abstract
We present the invariant structure of a Holomorphic Unified Field Theory in which gravity and gauge interactions arise from a single geometric framework. The theory is formulated on a product principal bundle, with one connection, and curvature equipped with a Hermitian field on a complexification of spacetime. From a single $\mathrm{Diff}(M)\times G$–invariant action, variation yields the Einstein and Yang–Mills equations together with their paired Bianchi identities. A compatibility condition is implemented either definitionally or through an auxiliary penalty functional. It enforces that the antisymmetric part of our Hermitian field is exactly the gauge field’s curvature on the real slice.
My methods are really methods of working and thinking; this is why they have crept in everywhere anonymously.
Emmy Noether, Letter to Helmut Hasse (1931)
1 Introduction
In 1945, Albert Einstein wrote, “Every attempt to establish a unified-field theory must start, in my opinion, from a group of transformations.” This means that unification is a symmetry-first program [1]. Emmy Noether had the view that any candidate must be built from a single invariant action [2]. From that one symmetry comes conserved currents and differential identities and or Bianchi-type relations [2, 3]. Noether used general relativity (GR) as the archetype for an infinite-dimensional symmetry diffeomorphisms, showing how conservation becomes a consequence of symmetry, not an add-on. The Holomorphic Unified Field Theory (HUFT) implements Einstein’s symmetry-first mandate by choosing one transformation structure and building geometry via a single metric-compatible connection plus a single gauge connection. It satisfies Noether by using one invariant action whose variations give both the Einstein and Yang–Mills equations. At the same time, the same symmetry yields Bianchi identities and conserved currents [4, 5, 6, 7]. HUFT solves the dynamical part of unification in its own sense with a single UV connection and coupling, regulated variational mechanism [8, 9, 10]. It does not yet solve the unique GUT-style unification unless we add and prove a geometric or topological locking principle that fixes the relative normalization and uniquely determines the internal group [11, 12].
We develop a symmetry–first formulation of HUFT, whose aim is to package gravity and gauge interactions into a single geometric structure without introducing new propagating degrees of freedom. The guiding principle is that dynamics should follow from one invariant action built from one set of geometric data on a single underlying bundle, so that the familiar Einstein and Yang–Mills sectors are recovered as two aspects of the same symmetry [5, 6]. In this sense, unification here means shared kinematics, identities, and variational origin is not a proliferation of fields or ad hoc couplings.
We work with a product principal bundle carrying a single connection that contains both the spin connection and the internal gauge connection [5]. We also introduce a Hermitian packaging of the spacetime metric with an antisymmetric two–form valued in the gauge algebra. Crucially, that antisymmetric part is not given independent dynamics, either by definition or through an algebraic compatibility term in the action, it is tied pointwise to the gauge curvature on the real slice of the theory [7]. All index operations and geometric constructions are performed with the ordinary Lorentzian metric [3]. This ensures that the holomorphic language functions as a bookkeeping device that makes the unification transparent while remaining mathematically safe.
We obtain from this setup two core results; first, at the classical level, the unified theory on the real slice is equivalent to the standard system of general relativity plus Yang–Mills with the same matter content [3, 4]. The conservation laws and Bianchi identities arise as a single Noether identity associated with diffeomorphism and internal gauge invariance [2, 7]. Second, at the level of formal path integrals, integrating out the auxiliary two–form reduces the functional measure to that of the standard theory, indicating that no additional propagating modes or counterterms are introduced by the holomorphic packaging [10, 13, 14]. We are explicit about the status of this second statement, as it follows established physics practice but is not presented as a measure–theoretic proof.
Beyond demonstrating equivalence, the paper advances two structural claims. The first is a minimality statement that if one insists on diffeomorphism– and gauge–natural, second–order dynamics built from the unified geometric data, the Einstein–Hilbert and Yang–Mills terms are singled out up to boundary and topological contributions [3, 6]. The second is an internal–group constraint, the holomorphic and Yang–Mills structure, together with chirality and charge integrality, points toward a unitary internal group with the minimal viable choice breaking to the Standard Model on the real slice [11, 12, 15].
While the mathematical structures we employ, Hermitian frame bundles, spin bundles, and holomorphic stability criteria, are well established in differential and complex geometry [16, 17], the novelty of this work lies in their assembly into a single unification framework. The complexified metric $g=h+iB$ is interpreted such that its antisymmetric part is constrained to equal the Yang-Mills curvature, thereby formulating gravitational and gauge data into one Hermitian structure. This identification, together with the formulation on the product bundle $P_{Spin(1,3)}\times_{M}P_{G}$ yields a single invariant action reproducing Einstein-Hilbert and Yang-Mills dynamics. The originality of the theory is not in introducing new mathematical objects, but in recombining standard geometric ingredients into a symmetry-driven unification principle. The Hermitian metric is constrained pointwise to be the Yang-MMills curvature, so that gravity and gauge interactions are implemented in a single Hermitian structure, while remaining dynamically equivalent to Einstein-Hilbert and plus Yang-Millss. In this sense, the work is in spirit symmetry based on Noetherian unification principles, employing a holomorphic bundle formalism.
Our goal is therefore twofold, to give a self–contained account of the invariant structure that makes the unification precise and mathematically consistent, and to delineate the exact sense in which the construction is equivalent to known physics while clarifying what is genuinely new and what is not. The remainder of the paper spells out the geometric data and symmetries, derives the field equations and identities from a single action, proves the equivalence at the classical level, analyses the formal quantum reduction, and discusses the minimality and internal–group considerations, together with open directions on anomaly data, matter–bundle building, and measure–theoretic rigour. We believe that this can and should be compared with past theories to induce a discussion about how a unified theory should be interpreted and formulated [18, 19, 20, 21, 22, 23, 24, 25].
2 Invariant Geometry of HUFT
We will show that the Hermitian metric as given by HUFT [26, 27, 28] does satisfy a unified theory as described by Albert Einstein [1] and Emmy Noether [2]. We let $M$ be an oriented, time–orientable, spin 44–manifold on the real slice. We let $P_{\mathrm{Spin}}\to M$ be the $\mathrm{Spin}(1,3)$ frame bundle and $P_{G}\to M$ a principal GG–bundle. We define the product principal bundle:
$P_{\mathrm{tot}}\;=\;P_{\mathrm{Spin}}\times_{M}P_{G},\qquad H:=\mathrm{Spin}(1,3)\times G,\qquad\mathfrak{h}=\mathfrak{spin}(1,3)\oplus\mathfrak{g}.$ (1)
Now we let $h\in\Gamma(S^{2}T^{*}M)$ be the smooth sections of the bundle of symmetric 2-forms with signature (+++)(-+++) be a Lorentzian metric on $M$ and $\mathcal{A}=(\omega,A)$ an HH–connection on $P_{\mathrm{tot}}$, with curvatures:
$\mathcal{F}\;=\;\mathrm{d}\mathcal{A}+\mathcal{A}\wedge\mathcal{A}\;=\;(R,F),$ (2)
where $R\in\Omega^{2}(M,\mathfrak{spin}(1,3))$ is the Riemann curvature 22–form of $\omega$ and $F\in\Omega^{2}(M,\mathrm{ad}P_{G})$ is the Yang–Mills field strength of $A$.
We now derive the internal group from the fiber geometry. We have $E\to M$ as the matter bundle of rank $n$ with Hermitian metric $h_{E}$, the first Chern class of the complex vector bundle is zero $c_{1}(E)=0$, the determinant line bundle $\text{det}E$ is topologically trivial so no net U(1) flux meaning the structure group can drop from $U(n)$ to $SU(n)$, a nowhere-vanishing holomorphic volume form $\Omega_{E}\in H^{0}(M,\wedge^{n}E^{*})$ meaning that there exists a global holomorphic section of the determinant line bundle $\wedge^{n}E^{*}\cong(\det E)^{*}$ which never vanishes on $M$ ion local holomorphic frames $\{e_{i}\}_{i=1}^{n}$ of $E$, the form can be written as $\Omega_{E}=f(z)\,e^{1}\wedge\cdots\wedge e^{n}$ with a holomorphic function $f(z)\neq 0$ everywhere. The existence of such a nowhere–vanishing section trivializes the determinant line bundle, implying $\det E\simeq\mathcal{O}_{M}$ and hence $c_{1}(E)=0$. Geometrically, $\Omega_{E}$ plays the role of a holomorphic volume element on the fibers of $E$, fixing an intrinsic orientation and allowing us to identify totally antisymmetric tensors via the holomorphic epsilon symbol. The subgroup of $GL(n,\mathbb{C})$ preserving both the Hermitian metric $h_{E}$ and $\Omega_{E}$ is $SU(n)$ and thus the data $(E,h_{E},\Omega_{E})$ define an $SU(n)$-structure on the bundle. Physically, this reduction eliminates the overall $U(1)$ determinant factor, enforcing traceless gauge transformations and ensuring charge quantization in the $SU(n)$ fiber, as required for unified models such as $SU(5)$. And we assume slope-stable so the compatible connection is Hermitian–Yang–Mills (HYM) [40, 41]. Preservation of $(h_{E},\Omega_{E})$ reduces the structure group of $E$ to $\mathrm{SU}(n)$, and the HYM connection has holonomy in $\mathrm{SU}(n)$. We let $Y\in\mathfrak{su}(n)$ be the traceless character compatible with $\Omega_{E}$. We normalize $Y$ by the calibrated pairing so that:
$\frac{1}{8\pi^{2}}\int_{M}\mathrm{Tr}(F_{Y}\wedge F_{Y})\in\mathbb{Z},$ (3)
fixing the abelian charge units without convention. Imposing a single adjoint breaking on the real slice, the hypercharge integrality above, and chiral, anomaly-free matter for one family from associated bundles of $E$, the minimal rank is $n=5$ [12, 15]. Therefore, the maximal compatible internal symmetry is:
$G=\mathrm{SU}(5)\ \longrightarrow\ \mathrm{SU}(3)_{c}\times\mathrm{SU}(2)_{L}\times\mathrm{U}(1)_{Y},$ (4)
under a holomorphic adjoint ($\mathbf{24}$) reduction, with the standard hypercharge generator [11, 12]. In what follows, we take $G=\mathrm{SU}(5)$ above the mass-energy scale $M_{*}$, with a holomorphic adjoint reduction to the Standard Model group on the real slice. Given the rank-$n$ holomorphic bundle $E\to M$ with Hermitian metric on the internal vector bundle $h_{E}$ and a nowhere-vanishing holomorphic volume form $\Omega_{E}\in H^{0}(M,\wedge^{n}E^{*})$, the structure group reduces as:
[…]
The preservation of $h_{E}$ picks unitary changes of frame, such as $\mathrm{U}(n)\subset\mathrm{GL}(n,\mathbb{C})$. Preservation of $\Omega_{E}$ forces $\det=1$ on $\mathrm{U}(n)$, giving $\mathrm{SU}(n)$. If, in addition, $c_{1}(E)=0$ and $E$ is slope-stable, the unique compatible connection is Hermitian–Yang–Mills and its holonomy lies in $\mathrm{SU}(n)$ [40, 41]. The internal gauge symmetry is the automorphism group preserving $(h_{E},\Omega_{E})$, namely $G=\mathrm{SU}(n)$. Imposing chirality and anomaly cancellation for one SM family, hypercharge integrality via $\frac{1}{8\pi^{2}}\int\mathrm{Tr}(F_{Y}\wedge F_{Y})\in\mathbb{Z}$, and a single adjoint breaking on the real slice fixes the minimal rank to $n=5$, we have:
$G=\mathrm{SU}(5)\ \longrightarrow\ \mathrm{SU}(3)_{c}\times\mathrm{SU}(2)_{L}\times\mathrm{U}(1)_{Y}.$ (6)
Now we take $\kappa$ to denote the Killing form on $\mathfrak{g}$ and $\langle\cdot,\cdot\rangle_{h}$, the fiberwise inner product induced by $h$ on differential forms. We say that gravity and gauge interactions are fully geometrically unified if there exists a single principal HH–bundle $P_{\mathrm{tot}}\to M$, a single connection $\mathcal{A}$ on $P_{\mathrm{tot}}$ with curvature $\mathcal{F}$, and a single $\mathrm{Diff}(M)\times H$–invariant action $S[h,\mathcal{A},\Psi]$ such that the Euler–Lagrange equations are equivalent to the Einstein equations for $h$ coupled to the Yang–Mills equations for $A$ and matter $\Psi$. The unique Bianchi identity $D_{\mathcal{A}}\mathcal{F}=0$ simultaneously yields the Riemann and Yang–Mills Bianchi identities [7, 3]. On $(M,P_{\mathrm{tot}},h,\mathcal{A})$ as above, we define:
$S[h,\mathcal{A},\Psi]=\int_{M}\sqrt{|h|}\left(\frac{1}{2\kappa_{\rm grav}}\,R(h)-\frac{1}{4}\,\langle F,F\rangle_{h}+\mathcal{L}_{\rm matter}(\Psi;h,\mathcal{A})\right),$ (7)
where $\langle F,F\rangle_{h}$ is computed using the Killing form $\kappa$ on g and the metric $h$. Then, We wouSS is invariant under $\mathrm{Diff}(M)\times H$. The unique Bianchi identity $D_{\mathcal{A}}\mathcal{F}=0$ splits as $D_{\omega}R=0$ and $D_{A}F=0$. Under an HH–gauge transformation $g:M\to H$:
$\mathcal{A}\mapsto g^{-1}\mathcal{A}g+g^{-1}\,\mathrm{d}g,\qquad\mathcal{F}\mapsto g^{-1}\mathcal{F}g,$ (8)
hence $F\mapsto g^{-1}Fg$ on the internal block. Because the Killing form $\kappa$ on $\mathfrak{g}$ is $\mathrm{Ad}$–invariant, $\mathrm{tr}_{\kappa}(F\wedge*_{h}F)$ is gauge invariant. Each integrand in $S$ is a scalar density built from $h$ (via $_{h}$ and contractions), so the action is also invariant under $\text{Diff}(M)$. The Euler–Lagrange equations of $S$ are:
$G_{\mu\nu}(h)$ $=\kappa_{\rm grav}\,\big(T^{\rm YM}_{\mu\nu}+T^{\rm matter}_{\mu\nu}\big),$ (9)
$D_{A}\left(*_{h}F\right)$ $=J_{\rm matter},$ (10)
the Einstein equations for $h$ coupled to the Yang–Mills equations for $\mathcal{A}$ with the same connection $A$ acting on matter. Hence the data $(P_{\mathrm{tot}},\mathcal{A},h)$ satisfy the Definition of full geometric unification.
The Einstein–Hilbert term and $\langle F,F\rangle_{h}$ are $\mathrm{Diff}(M)$–invariant by construction as each integrand is a scalar density obtained from h and F using only natural tensor operations, so the whole action is invariant under pullbacks by diffeomorphisms. Gauge invariance under $H$ holds because $F\mapsto g^{-1}Fg$ and $\kappa$ are $\mathrm{Ad}$–invariant, so $\mathrm{tr}_{\kappa}(F\wedge*_{h}F)$ is gauge invariant. The matter Lagrangian is assumed to be constructed from $h$ and the covariant derivative $D_{\mu}=\partial_{\mu}+\frac{1}{4}\omega^{ab}{}_{\mu}\gamma_{ab}+i\,g\,A_{\mu}^{A}T_{A}$, meaning it is also $\mathrm{Diff}(M)\times H$–invariant this is so that every derivative of a field is replaced by one that transforms covariantly under both spacetime and internal symmetries. Since $\mathfrak{h}=\mathfrak{spin}(1,3)\oplus\mathfrak{g}$, the covariant derivative $D_{\mathcal{A}}$ splits as $D_{\mathcal{A}}=(D_{\omega},D_{A})$ on the two summands. Therefore $(D_{\omega}R,\;D_{A}F)=D_{\mathcal{A}}\mathcal{F}=0$, the Riemann and Yang–Mills Bianchi identities simultaneously. Varying $S$ with respect to $A$ and integrating by parts yields:
$\delta_{A}S\;=\;-\frac{1}{2}\int_{M}\sqrt{|h|}\,\langle D_{A}(*_{h}F)-J_{\rm matter},\,\delta A\rangle\quad\Rightarrow\quad D_{A}(*_{h}F)=J_{\rm matter},$ (11)
the Yang–Mills equations (10). Varying $S$ with respect to $h$ gives:
$\delta_{h}S\;=\;\frac{1}{2}\int_{M}\sqrt{|h|}\,\big(G_{\mu\nu}-\kappa_{\rm grav}\,T_{\mu\nu}\big)\,\delta h^{\mu\nu},$ (12)
where $T_{\mu\nu}=T^{\rm YM}_{\mu\nu}+T^{\rm matter}_{\mu\nu}$ and $T^{\rm YM}_{\mu\nu}=\mathrm{tr}_{\kappa}\left(F_{\mu\alpha}F_{\nu}{}^{\alpha}-\frac{1}{4}h_{\mu\nu}F_{\alpha\beta}F^{\alpha\beta}\right)$, this implies (9). Since both sectors arise from one variational principle on the same $(h,A)$, the Definition is satisfied.
We let $M_{\mathbb{C}}$ be a complexification of $M$ with complex coordinates $z^{\mu}=x^{\mu}+iy^{\mu}$, and let $g=h+iB$ be a Hermitian tensor field on $M_{\mathbb{C}}$ with $h\in\Gamma(S^{2}T^{*}M)$ and $B\in\Omega^{2}(M,\mathrm{ad}P_{G})$ pulled back from $M$. Now we impose the metric-compatibility condition with the same master connection:
$\nabla_{A}g=0\quad\text{on }M_{\mathbb{C}}\,$, (13)
and restrict to the real slice $y=0$. The real and imaginary parts of (13) give:
$\nabla^{\rm LC}_{\mu}h_{\alpha\beta}=0,\qquad D_{A}B=0\quad\text{on }M,$ (14)
so $h$ is the spacetime metric, and $B$ is an $\mathrm{ad}P_{G}$–valued closed 22–form under $D_{A}$. In the dynamical theory, either by definition or by adding a holomorphic penalty term $\int\sqrt{|h|}\,\langle B-F,B-F\rangle_{h}$ and using the Euler–Lagrange equation, we obtain $B=F$ on the real slice [7, 3], we augment the action by the holomorphic penalty:
$S_{\text{pen}}[h,A,B]=\frac{\lambda}{2}\int_{M}\sqrt{|h|}\,\langle B-F,\;B-F\rangle_{h}=\frac{\lambda}{2}\int_{M}\kappa\big((B-F)\wedge*_{h}(B-F)\big),$ (15)
where $F=dA+A\wedge A$, $\kappa$ is the Killing form on $\mathfrak{g}$, and $_{h}$ is the Hodge operator determined by $h$. We treat $h,A,B$ as independent fields and we work on an oriented $M$, drop the boundary terms, and use that $_{h}$ is an isomorphism on $2$-forms in four dimensions. Since $S_{\text{pen}}$ is algebraic in $B$:
$\delta_{B}S_{\text{pen}}=\lambda\int_{M}\kappa\big(\delta B\wedge*_{h}(B-F)\big)=\lambda\int_{M}\sqrt{|h|}\,\langle B-F,\;\delta B\rangle_{h}.$ (16)
Because $\delta B$ is arbitrary, the Euler–Lagrange equation from $\delta_{B}S_{\text{pen}}=0$ is:
$_{h}(B-F)=0\quad\Longrightarrow\quad B-F=0.$ (17)
Using $\delta F=D_{A}\delta A$ and integrating by parts covariantly:
$\delta_{A}S_{\text{pen}}$ $=-\lambda\int_{M}\kappa\big((B-F)\wedge*_{h}D_{A}\delta A\big)=\lambda\int_{M}\kappa\big(D_{A}(*_{h}(B-F))\wedge\delta A\big),$ (18)
so the AA–equation of motion contributed by $S_{\text{pen}}$ is:
$D_{A}\big(*_{h}(B-F)\big)=0.$ (19)
Together with (17), this is automatically satisfied. On the real slice, we also have $D_{A}B=0$ and the Bianchi identity $D_{A}F=0$, hence $D_{A}(B-F)=0$ and $D_{A}\big(*_{h}(B-F)\big)=0$.
This means the Euler–Lagrange equation from varying $B$ in (15) enforces the pointwise identification $B=F\quad\text{on }M,$ so the antisymmetric sector coincides with the Yang–Mills curvature on the real slice. The antisymmetric sector of $g$ is the gauge curvature, while all index operations use $h$. Because $S$ is $\mathrm{Diff}(M)\times H$–invariant and built from $(h,A)$, the single Noether identity yields both covariant conservations:
$\nabla^{h}_{\mu}T^{\mu}{}_{\nu}=0,\qquad D_{A}J_{\rm matter}=0,$ (20)
and matter couples by the same covariant derivative $D_{\mu}$ that contains both the spin and internal connections. This shows parallel transport, curvature, symmetries, and dynamics are all governed by the single geometric object $\mathcal{A}$ on $P_{\mathrm{tot}}$, with $h$ supplying measurement such as Hodge duals, index operations. This is precisely full geometric unification in the sense of Definition 1, achieved without inverting the full Hermitian field $g=h+iB$.
Figure 1: The top ribbon is the Spin frame bundle $\pi_{\mathrm{Spin}}:P_{\mathrm{Spin}}\to M$ carrying the tetrad or coframe and spin connection $(e^{\alpha},\omega)$; the bottom ribbon is the internal gauge bundle $\pi_{G}:P_{G}\to M$ carrying the gauge potential and curvature $(A,F)$. The dashed oval indicates the fiber product over a common base point $x\in M$.
In this s is the frame choice on $P_{Spin}$ as $s_{Spin}:M\to P_{Spin}$ picks an orthonormal frame. Pulling back the canonical forms on the bundle gives the coframe and the local Spin connection on M:
$e^{\alpha}=s_{Spin}^{*}(\vartheta^{\alpha}),\quad\omega=s_{Spin}^{*}(\Omega),$ (21)
where $\Omega$ is the principal connection on $P_{Spin}.$
The bundle geometry of the bundle $P_{Spin}\times_{M}P_{G}$ as connected by $\mathcal{F,A}$, we have $e^{\alpha}$ as the coframe one form that converts spacetime indices ($\mu,\nu,$...) to local Lorentz indices ($\alpha,\beta,$...). We have a set of one forms on $M$:
$e^{\alpha}=e_{\mu}^{\alpha}dx^{\mu},\quad\alpha=0,1,2,3.$ (22)
The dual to the frame vector fields $e_{\alpha}=e^{\mu}_{\alpha}\partial_{\mu}$ with:
$e^{\alpha}(e_{\beta})=\delta_{\beta}^{\alpha},\quad e^{\mu}_{\alpha}e^{\beta}_{\mu}=\delta_{\alpha}^{\beta},\quad e^{\alpha}_{\mu}e^{\nu}_{\alpha}=\delta_{\mu}^{\nu}.$ (23)
This builds the spacetime metric from the Minkowski metric $\eta_{\alpha\beta}=\text{diag}(-+++)$:
$h_{\mu\nu}=\eta_{\alpha\beta}e_{\mu}^{\alpha}e^{\beta}_{\nu},\quad h^{\mu\nu}=\eta^{\alpha\beta}e^{\mu}_{\alpha}e_{\beta}^{\nu},\quad\sqrt{|h|}=\text{det}(e^{\alpha}_{\mu}).$ (24)
In this $e^{\alpha}$ provides the soldering between the Spin bundle and the the tangent bundle (TM) of the manifold $M$.
For each point $x\in M$, the tangent space $T_{x}M$ is the vector space of velocities of smooth curves through $x$ or equivalently, derivations on $C^{\infty}(M)$ at $x$. The tangent bundle is the disjoint union of all tangent spaces, with a smooth bundle structure:
$TM\;=\;\bigsqcup_{x\in M}T_{x}M,\qquad\pi_{TM}:TM\to M,\quad(x,v)\mapsto x.$ (25)
If $\dim M=n$, here $n=4$, then each fiber $T_{x}M\cong\mathbb{R}^{n}$. In a chart $(x^{\mu})$ on $M$, a tangent vector is $v=v^{\mu}\partial_{\mu}\!\mid_{x}$, and a point of $TM$ has coordinates $(x^{\mu},v^{\mu})$. We have a vector field, which is a section of $TM$:
$X:M\to TM,\qquad\pi_{TM}\circ X=\mathrm{id}_{M},$ (26)
with components $X=X^{\mu}(x)\,\partial_{\mu}$. The cotangent bundle $T^{*}M$ has fibers of 1-forms so, the tetrad or coframe $e^{\alpha}=e^{\alpha}{}_{\mu}\,dx^{\mu}$ lives in $\Omega^{1}(M)=\Gamma(T^{*}M)$. The metric $h$ is a section of $S^{2}T^{*}M$ and identifies vectors or covectors through the musical isomorphisms  [^1] Musical isomorphisms have the form: $\sharp=\hat{g}:T^{*}M\to TM,\quad\flat=\hat{g}:TM\to T^{*}M,$ (27) these may as well be referred to as mutually inverse isomorphisms, and the (co)vectors obtained from this way are called metrically equivalent. [42, 43]:
$X^{\flat}=h_{\mu\nu}X^{\mu}dx^{\nu},\qquad\alpha^{\sharp}=h^{\mu\nu}\alpha_{\nu}\,\partial_{\mu}.$ (28)
We pick a local orthonormal basis $e_{\alpha}=e_{\alpha}{}^{\mu}\partial_{\mu}$ of $TM$. The coframe $e^{\alpha}=e^{\alpha}{}_{\mu}dx^{\mu}$ is its dual, and:
$h_{\mu\nu}=\eta_{\alpha\beta}\,e^{\alpha}{}_{\mu}\,e^{\beta}{}_{\nu}$. (29)
This is what we mean by the tetrad soldering $P_{\mathrm{Spin}}$ to $TM$: pulling back the solder form on the spin bundle gives $e^{\alpha}$, a $T^{*}M$-valued object.
For $\pi:M_{\mathbb{C}}\to M$ with $z^{\mu}=x^{\mu}+i\,y^{\mu}$, the tangent bundle of the complexification is $TM_{\mathbb{C}}$. An Ehresmann connection on $\pi$ gives, near $y=0$, a splitting:
$TM_{\mathbb{C}}\ \simeq\ \mathrm{Hor}\ \oplus\ \mathrm{Ver},\qquad\mathrm{Ver}=\ker(d\pi),$ (30)
so horizontal directions project to $TM$ while vertical directions are along the $y$-fibers.
If $s:M\to P_{Spin}$ is a local selection and $\vartheta^{\alpha}$ is the solder form, then $\vartheta^{\alpha}$ is the $\mathbb{R}^{1,3}$-valued one form on the orthonormal frame or Spin bundle that identifies tangent directions on $M$ with components in a chosen local Lorentz frame. It is the geometric device that connects the principal frame bundle to the base manifold’s tangent bundle.
We let $FM$ be the frame bundle of an $n$-manifold $M$, we have a point $u\in FM$ is a linear isomorphism $u:\mathbb{R}^{n}\to T_{x}M$, a frame at $x=\pi(u)$. The solder form $\vartheta\in\omega^{1}(FM);\mathbb{R}^{n}$ is defined by:
$\vartheta_{u}(X):=u^{-1}(d\pi_{u}(X))\in\mathbb{R}^{n},\quad X\in T_{u}FM.$ (31)
So this means, project a tangent vector $X$ at $u$ down to $T_{x}M$ through $d\pi$, then express it in the frame $u$ by applying $u^{-1}$. It has the property that if $X$ is tangent to the fiber, $d\pi(X)=0$, the $\vartheta(X)=0$. For the right action $R_{g}:FM\to FM,\quad g\in GL(n)$:
$(R_{g})^{*}\vartheta=g^{-1}\vartheta.$ (32)
On the orthonormal frame bundle $P_{SO(1,3)},$ this holds with $g\in SO(1,3)$.
For frames to the coframe on $e^{\alpha}$ on M, we chose a local section, $s:U\subset M\implies P_{SO(1,3)}$. We pull back the solder form:
$e^{\alpha}=s^{*}u^{\alpha}\in\Omega^{1}(U).$ (33)
These are the tetrad or coframe 1-forms (24).
The relation to the spin bundle is shown when we let $\lambda:P_{Spin}\implies P_{SO(1,3)}$ be the 2-to-1 form covering, where $\lambda$ is the bundle map that realizes the spin structure—the double cover from the Spin principal bundle to the orthonormal frame bundle. We pull $\vartheta$ back along $\lambda$ to get the spin bundle version:
$\vartheta_{Spin}=\lambda^{*}_{Spin}\vartheta^{\alpha}.$ (34)
This transforms as a Lorentz vector under Spin(1,3) rotations:
$e^{\alpha}\to\Lambda_{\beta}^{\alpha}(x)e^{\beta}.$ (35)
Given the Spin connection $\omega_{\beta}^{\alpha}$ on $P_{Spin}$, the Cartan structure quations are:
$T^{\alpha}=de^{\alpha}+\omega^{\beta}_{\alpha}\wedge e^{\beta},\quad R_{\beta}^{\alpha}=d\omega_{\gamma}^{\alpha}+\omega_{\gamma}^{\alpha}\wedge\omega_{\beta}^{\gamma},$ (36)
for Levi-Civita geometry $T^{\alpha}=0$.
[…]
We have the metric identities:
$\delta\sqrt{|h|}$ $=-\tfrac{1}{2}\sqrt{|h|}\,h_{\mu\nu}\,\delta h^{\mu\nu},$ (37)
$\delta R$ $=R_{\mu\nu}\,\delta h^{\mu\nu}+\nabla_{\mu}\!\big(\nabla_{\nu}\,\delta h^{\mu\nu}-\nabla^{\mu}\delta h\big),\quad\delta h:=h_{\alpha\beta}\delta h^{\alpha\beta}.$ (38)
therefore:
$\delta\!\big(\sqrt{|h|}R\big)=\sqrt{|h|}\,G_{\mu\nu}\,\delta h^{\mu\nu}+\text{(total derivative)}\!,\qquad G_{\mu\nu}:=R_{\mu\nu}-\tfrac{1}{2}h_{\mu\nu}R.$ (39)
For the gravitational part $S_{\rm EH}=\tfrac{1}{2\kappa}\int_{M}\!\sqrt{|h|}\,R$:
$\delta S_{\rm EH}=\frac{1}{2\kappa}\int_{M}\!\sqrt{|h|}\,G_{\mu\nu}\,\delta h^{\mu\nu}+\frac{1}{2\kappa}\int_{\partial M}\!\cdots$ (40)
The boundary term is cancelled by the Gibbons–Hawking–York term [30, 31]:
$S_{\rm GHY}=\frac{1}{\kappa}\int_{\partial M}\!\sqrt{|{\gamma}|}\,K,$ (41)
for Dirichlet data on $h$. For matter:
$S_{\rm m}=\int_{M}\!\sqrt{|h|}\,\mathcal{L}_{\rm matter}$ (42)
we define:
$T_{\mu\nu}:=-\frac{2}{\sqrt{|h|}}\frac{\delta S_{\rm m}}{\delta h^{\mu\nu}}.$ (43)
Stationarity $\delta S_{\rm EH}+\delta S_{\rm m}=0$ for arbitrary $\delta h^{\mu\nu}$ gives:
$G_{\mu\nu}(h)=\kappa\,T_{\mu\nu}\;.$ (44)
We write the YM action in differential-form notation:
$S_{\rm YM}=-\tfrac{1}{2}\int_{M}\!\langle F\wedge{*}_{h}F\rangle\;=\;-\tfrac{1}{4}\int_{M}\!\sqrt{|h|}\,\kappa_{AB}\,F^{A}_{\mu\nu}F^{B\,\mu\nu}.$ (45)
Since $\delta F=D_{A}(\delta A)$, we have:
$\delta S_{\rm YM}$ $=-\int_{M}\!\langle D_{A}\delta A\wedge{*}F\rangle\;=\;-\int_{M}\!d\langle\delta A\wedge{*}F\rangle+\int_{M}\!\langle\delta A\wedge D_{A}{*}F\rangle.$ (46)
With $\delta A|_{\partial M}=0$ or by adding the natural boundary term, the first term drops. Matter coupling defines the gauge current by:
$\delta S_{\rm m}\big|_{A}=\int_{M}\!\langle\delta A\wedge{*}J\rangle\quad\Longleftrightarrow\quad\frac{\delta S_{\rm m}}{\delta A_{\mu}^{A}}=\sqrt{|h|}\,J^{A\,\mu}.$ (47)
Stationarity $\delta(S_{\rm YM}+S_{\rm m})=0$ for arbitrary $\delta A$ yields the YM equations:
$D_{A}({*}_{h}F)={*}_{h}J\;\qquad\text{or in components }D_{\mu}F^{A\,\mu\nu}=J^{A\,\nu}\text{}.$ (48)
Varying $S_{\rm YM}$ with respect to $h^{\mu\nu}$ (using $\delta(\sqrt{|h|}F^{2})=\sqrt{|h|}\,[\,-\tfrac{1}{2}h_{\mu\nu}F^{2}+2\,F_{\mu\rho}F_{\nu}{}^{\rho}\,]\,\delta h^{\mu\nu}$) gives:
$T^{\rm YM}_{\mu\nu}=\kappa_{AB}\!\left(F^{A}_{\mu\rho}F^{B}{}_{\nu}{}^{\rho}-\tfrac{1}{4}h_{\mu\nu}F^{A}_{\rho\sigma}F^{B\,\rho\sigma}\right).$ (49)
Diffeomorphism invariance implies $\nabla^{\mu}G_{\mu\nu}\equiv 0$, contracted Bianchi implies $\nabla_{\mu}T^{\mu\nu}=0$ on shell. Gauge invariance implies $D_{A}D_{A}({*}F)\equiv 0$ implies covariant current conservation $D_{A}J=0$.
In the Palatini or tetrad formalism [32, 33, 34] one varies $(e^{a},\omega^{ab})$ independently, the torsionless solution reproduces the Levi–Civita case. If $M$ has a boundary, use $S_{\rm GHY}$ and the YM boundary term $\int_{\partial M}\!\langle\delta A\wedge{*}F\rangle$ to enforce Dirichlet data.
3 Real Slice Geometry
We will now explain real-slice geometry as a finer bundle, to show how the real slice geometry encodes symmetries in HUFT. By a finer bundle over $M$ we mean an enlargement of a given bundle $q:P\to M$ by adjoining an extra fiber $E_{y}\to M$, encoding the imaginary directions $y$ from the complexification $M_{\mathbb{C}}$:
$\widehat{P}\;:=\;P\times_{M}E_{y}\;\longrightarrow\;M,\qquad\pi_{1}:\widehat{P}\to P\text{ over }\mathrm{id}_{M},$ (50)
so that fiberwise $\widehat{P}_{x}\cong P_{x}\times(E_{y})_{x}$ and $\pi_{1}$ simply forgets $y$. In our setting $P=P_{\mathrm{tot}}=P_{\mathrm{Spin}}\times_{M}P_{G}$, so $\widehat{P}=(P_{\mathrm{Spin}}\times_{M}P_{G})\times_{M}E_{y}$. This is not a reduction of the structure group nor a refinement of an atlas, but it is an enrichment of the fiber used to keep track of the $y$-directions while all fields and dynamics are evaluated on the real slice $M$. We let $M$ be a smooth oriented, time–orientable spin 44–manifold, the real slice. We have $M_{\mathbb{C}}$ as a complexification with coordinates $z^{\mu}=x^{\mu}+iy^{\mu}$ and projection $\pi:M_{\mathbb{C}}\to M$, $\pi(x,y)=x$. The fibers $Y_{x}:=\pi^{-1}(x)$ are real 44–planes isomorphic to $\mathbb{R}^{4}$, defining a rank–$4$ real vector bundle $E_{y}\to M$ with total space canonically identified near $y=0$ with a neighborhood of $M\subset M_{\mathbb{C}}$.
We have $P_{Spin}\to M$ as the $Spin(1,3)$ frame bundle and $P_{G}\to M$ a principal GG–bundle with $G=SU(3)\times SU(2)\times U(1)$. Now we define the product bundle $P_{tot}:=P_{Spin}\times_{M}P_{G}$ with structure group $H:=Spin(1,3)\times G$. We refine this by adjoining the yy–fiber:
$\widehat{P}\;:=\;P_{tot}\times_{M}E_{y}\;\;\longrightarrow\;\;M,$ (51)
whose points encode local Lorentz frames, internal gauge frames, and the imaginary directions $y$ of the complexified manifold. Now we choose an HH–connection $\mathcal{A}=(\omega,A)$ on $P_{\mathrm{tot}}$ and an Ehresmann connection on $\pi:M_{\mathbb{C}}\to M$ that splits $TM_{\mathbb{C}}\simeq\mathsf{Hor}\oplus\mathsf{Ver}$ near $y=0$. Now we write the unified curvature:
$\mathcal{F}\;=\;\mathrm{d}\mathcal{A}+\mathcal{A}\wedge\mathcal{A}\;=\;R\oplus F,\qquad R\in\Omega^{2}(M,\mathfrak{spin}(1,3)),\;\;F\in\Omega^{2}(M,\mathrm{ad}P_{G}),$ (52)
and let $h\in\Gamma(\text{Sym}^{2}T^{*}M)$ be a Lorentzian metric, used for index operations and Hodge duals. On $M_{\mathbb{C}}$ we consider a Hermitian field
$g\;=\;h\;+\;i\,B,\qquad B\in\Omega^{2}\big(M,\mathrm{ad}P_{G}\big),$ (53)
and identify on the real slice $B=F$ either by definition or via a holomorphic compatibility term enforcing BF=0B-F=0 in the equations of motion. The base metric $h$ transforms under $\text{Diff}(M)$ as its Levi–Civita and spin connection $\omega$ is the $Spin(1,3)$ part of $A$. The curvature $R$ is the gravitational field strength; all contractions, Hodge duals, and stress–energy are built using $h$. The internal symmetry $G=SU(3)\times SU(2)\times U(1)$ acts on $P_{G}$; the unified gauge potential $A=A^{(3)}\oplus A^{(2)}\oplus A^{(1)}$ and field strength $F=F^{(3)}\oplus F^{(2)}\oplus F^{(1)}$ live in $\Omega^{2}(M,\mathrm{ad}P_{G})$. On the real slice, the antisymmetric part of the Hermitian metric equals this curvature: $g_{[\mu\nu]}\equiv B_{\mu\nu}=F_{\mu\nu}$. Finer bundle with $y$ inside, The yy–fiber $E_{y}$ records the imaginary directions of $M_{\mathbb{C}}$. The Ehresmann splitting ties variations in $y$ to horizontal transport on $M$. In the holomorphic description, the mixed (x,y)(x,y)–geometry packages the gauge sector into the antisymmetric two–form $B$ while preserving standard spacetime geometry in $h$. By Noether I & II the single invariant action: A single $\text{Diff}(M)\times G$–invariant action:
$S[h,\mathcal{A},\Psi]=\int_{M}\sqrt{|h|}\left(\frac{1}{2\kappa}R(h)-\frac{1}{4}\langle F,F\rangle_{h}+\mathcal{L}_{\rm matter}(\Psi;h,\mathcal{A})\right),$ (54)
yields, by variation:
$G_{\mu\nu}(h)=\kappa\,T_{\mu\nu}(F,\Psi;h),\qquad D_{A}(*_{h}F)=J_{\rm matter},$ (55)
and by Noether’s second theorem the unified Bianchi identities $D_{\omega}R=0$, $D_{A}F=0$, together with covariant current conservation $\nabla_{\mu}T^{\mu\nu}=0$, $D_{A}J=0$. The same symmetry explains the conservation laws. Writing $G=SU(3)\times SU(2)\times U(1)$ with Lie algebra $\mathfrak{g}=\mathfrak{su}(3)\oplus\mathfrak{su}(2)\oplus\mathfrak{u}(1)$ and Killing form $\kappa$, we decompose:
$A=A_{\mu}dx^{\mu}=A^{(3)a}_{\mu}T^{(3)}_{a}\,dx^{\mu}\oplus A^{(2)i}_{\mu}T^{(2)}_{i}\,dx^{\mu}\oplus A^{(1)}_{\mu}Y\,dx^{\mu},$ (56)
$F=dA+A\wedge A=F^{(3)}\oplus F^{(2)}\oplus F^{(1)},\qquad\langle F,F\rangle_{h}=\kappa_{AB}F^{A}_{\mu\nu}F^{B\,\mu\nu}.$ (57)
On the real slice $y=0$:
$g_{[\mu\nu]}\;\equiv\;B_{\mu\nu}\;=\;F_{\mu\nu}^{(3)}\oplus F_{\mu\nu}^{(2)}\oplus F_{\mu\nu}^{(1)}\;,$ (58)
and all index operations use only $h^{\mu\nu}$. Thus, the antisymmetric sector encodes the entire SM gauge curvature, while the symmetric sector encodes spacetime geometry.
The refined bundle $\widehat{P}$ packages diffeomorphisms through $h,\omega$ and SM gauge symmetries via $A,F$ and therefore $g_{[\mu\nu]}$, all governed by one symmetry group $\text{Diff}(M)\times G$ and one invariant action. This realizes Einstein’s symmetry-first mandate and Noether’s criterion that conservation and identities follow from the same symmetry.
Figure 2: The upper ribbon is $M_{\mathbb{C}}$ with coordinates $z^{\mu}=x^{\mu}+i\,y^{\mu}$, the lower ribbon is the real slice $M$. The projection $\pi:M_{\mathbb{C}}\!\to M$ sends $(x,y)\mapsto x$. The black dots mark the point $s_{0}(x)=(x,0)\in M_{\mathbb{C}}$ (label $s$) and its projection $x\in M$ (label $\pi$), so that $\pi\circ s_{0}=\mathrm{id}_{M}$. The label $\mathbb{P}_{y}$ denotes the rank–$4$ real vector bundle $E_{y}\to M$ of imaginary directions: for each $x\in M$, the fiber $Y_{x}=\pi^{-1}(x)\cong\mathbb{R}^{4}$ carries coordinates $y^{\mu}$.
The diagram shows the complexification $\pi:M_{\mathbb{C}}\!\to M$, with the upper ribbon $M_{\mathbb{C}}$ and the lower ribbon the real slice $M$. The point labelled $s$ is the image of the zero section $s_{0}:M\hookrightarrow M_{\mathbb{C}}$, $s_{0}(x)=(x,0)$; the lower point labelled $\pi$ is the basepoint $x\in M$, so $\pi\!\circ s_{0}=\mathrm{id}_{M}$. The curved line between them indicates the fiber $Y_{x}:=\pi^{-1}(x)$ over $x$, which is a real 44–plane isomorphic to $\mathbb{R}^{4}$ and coordinatized by the imaginary directions $y^{\mu}$. We denote by $\mathbb{P}_{y}$ in the figure our $E_{y}$ in the text the rank–$4$ real vector bundle $E_{y}\to M$ whose fiber at $x$ is $Y_{x}$; near $y=0$ its total space is canonically identified with a neighborhood of $M\subset M_{\mathbb{C}}$. This bookkeeping of the yy–directions, lets us speak of a refined finer bundle over $M$ that remembers the imaginary displacements while all fields are ultimately evaluated on the real slice.
4 Symmetry Completeness and its Physical Equivalence
As a sanity check, we will show that we recover symmetry completeness and explore its physical equivalence. We have $M$ as an oriented, time-orientable spin 44–manifold and let $P_{\text{Spin}}\to M$ be the $\text{Spin}(1,3)$ frame bundle and $P_{G}\to M$ a principal GG–bundle with $G=SU(3)\times SU(2)\times U(1)$. Set $P_{\text{tot}}=P_{\text{Spin}}\times_{M}P_{G}$ with structure group $H=\text{Spin}(1,3)\times G$. Now let $h\in\Gamma(\text{Sym}^{2}T^{*}M)$ be a Lorentzian metric and $\mathcal{A}=(\omega,A)$ an HH–connection with curvature $\mathcal{F}=\mathrm{d}\mathcal{A}+\mathcal{A}\wedge\mathcal{A}=(R,F)$, where $R$ is the Riemann curvature $2$-form and $F$ the Yang–Mills field strength. On the complexified bundle of $(0,2)$-tensors, define the Hermitian field as in Eq( 53) and on the real slice impose the compatibility $B=F$ either as a definition or via a holomorphic penalty term enforcing BF=0B-F=0 in the equations of motion.
We consider the single $\text{Diff}(M)\times G$–invariant action:
$S[h,\mathcal{A},\Psi]=\int_{M}\sqrt{|h|}\,\Big(\tfrac{1}{2\kappa}R(h)-\tfrac{1}{4}\,\langle F,F\rangle_{h}+\mathcal{L}_{\rm matter}(\Psi;h,\mathcal{A})\Big).$ (59)
where $\langle\cdot,\cdot\rangle_{h}$ is the fiberwise inner product induced by $h$ and the Killing form on $\mathfrak{g}$, and $\Psi$ denotes matter fields. Here $\mathcal{A}=(\omega,A)$ with curvature $\mathcal{F}=\mathrm{d}\mathcal{A}+\mathcal{A}\wedge\mathcal{A}=(R,F)$, so $D_{\mathcal{A}}=(D_{\omega},D_{A})$ on $\mathfrak{spin}(1,3)\oplus\mathfrak{g}$.
Theorem 1
For the data above, the following hold on the real slice $B=F$:
  1. (i)
    $S$ is invariant under $\text{Diff}(M)\times H$.
  2. (ii)
    Noether II for $\text{Diff}(M)\times G$ yields the unified Bianchi identities $D_{\omega}R=0$ and $D_{A}F=0$; Noether I yields the covariant conservations $\nabla_{\mu}T^{\mu\nu}=0$ and $D_{A}J=0$.
  3. (iii)
    The Euler–Lagrange equations of $S$ are:
    $G_{\mu\nu}(h)=\kappa\big(T^{\rm YM}_{\mu\nu}+T^{\rm matter}_{\mu\nu}\big),\qquad D_{A}(*_{h}F)=J_{\rm matter},$ (60)
    with the standard Yang–Mills stress tensor $T^{\rm YM}_{\mu\nu}={\rm tr}\left(F_{\mu\alpha}F_{\nu}{}^{\alpha}-\tfrac{1}{4}h_{\mu\nu}F_{\alpha\beta}F^{\alpha\beta}\right)$ and gauge current $J$ defined from $\mathcal{L}_{\rm matter}$ by minimal coupling.
  4. (iv)
    (Classical equivalence of physics) The map:
    $\Phi:\ (h,A,\Psi)\ \longmapsto\ (g=h+iF,\ \mathcal{A}=(\omega(h),A),\ \Psi),$ (61)
    induces a bijection between solution spaces modulo $\text{Diff}(M)\times G$:
    $\mathfrak{Sol}_{\rm EYM+matter}/(\text{Diff}\times G)\;\cong\;\mathfrak{Sol}_{\rm unified\ real\ slice}/(\text{Diff}\times G).$ (62)
    Hence, all classical observables and their conservation laws coincide.
  5. (v)
    (Quantum equivalence, formal) If one includes a gauge-invariant penalty $\frac{\lambda}{2}\int\sqrt{|h|}\,\langle B-F,B-F\rangle_{h}$ and integrates out $B$, then, for $\lambda\to\infty$, the generating functional reduces to that of the standard EYM++matter theory, so perturbative correlators and SS–matrix elements agree.
Under HH–gauge transformations $F\mapsto g^{-1}Fg$ and the Killing form is Ad–invariant, so $\mathrm{tr}(F\wedge*_{h}F)$ is gauge invariant. Each term in (59) is a scalar density, hence $\text{Diff}(M)$–invariant. For a connection $A$ on a principal HH–bundle, $D_{A}F=0$ is the unified Bianchi identity, which splits as $D_{\omega}R=0$ and $D_{A}F=0$. Diffeomorphism invariance implies the contracted Bianchi identity $\nabla_{\mu}G^{\mu\nu}=0$ and hence $\nabla_{\mu}T^{\mu\nu}=0$ on shell GG–invariance implies $D_{A}J=0$. Varying $A$ with $h$ fixed gives:
$\delta S_{\rm YM}=-\tfrac{1}{2}\int\sqrt{|h|}\,\langle D_{A}(*_{h}F),\delta A\rangle,\qquad\delta S_{\rm matter}=\tfrac{1}{2}\int\sqrt{|h|}\,\langle J_{\rm matter},\delta A\rangle,$ (63)
so $D_{A}(*_{h}F)=J_{\rm matter}$. Varying $h$ yields:
$\delta S=\tfrac{1}{2}\int\sqrt{|h|}\,\big(G_{\mu\nu}-\kappa\,T_{\mu\nu}\big)\,\delta h^{\mu\nu},$ (64)
with $T_{\mu\nu}=T^{\rm YM}_{\mu\nu}+T^{\rm matter}_{\mu\nu}$, giving the Einstein equation.
Given any EYM++matter solution $(h,A,\Psi)$, define $(g,A,\Psi)$ by $g=h+iF$ and $\mathcal{A}=(\omega(h),A)$. Then $(g,A,\Psi)$ solves the unified Euler–Lagrange system on the real slice because (60) hold, and $B=F$ by construction. Any unified real-slice solution $(g,A,\Psi)$ has $g=h+iB$ with $B=F$ from the BB–equation, and its $(h,A,\Psi)$ satisfies (60). Both constructions are natural with respect to $\text{Diff}(M)\times G$, so they descend to a bijection of quotient solution spaces. Classical observables built from $(h,A,\Psi)$, such as fluxes, charges, stress-energy, coincide with those built from $(g,A,\Psi)$ restricted to the real slice.
We add $S_{\lambda}[B,A;h]=\tfrac{\lambda}{2}\int\sqrt{|h|}\,\langle B-F,B-F\rangle_{h}$. The BB–integral is Gaussian:
$\int DB\,e^{\,iS_{\lambda}[B,A;h]}\;\propto\;(\det\lambda)^{-1/2}\,e^{\,i\cdot 0}\,\delta[B-F]\ \xrightarrow{\ \lambda\to\infty\ }\ \delta[B-F].$ (65)
Thus, the unified partition function reduces to the standard EYM++matter one up to an overall constant, with identical gauge-fixing or ghost structure. Gauge and diffeo invariance and BRST remain as in the standard theory.
The consequences of this are that we have the same symmetries; $\text{Diff}(M)\times G$ invariance, Bianchi identities, and covariant current and stress-energy conservation are identical to GR++YM. The same observables, such as charges, fluxes, and classical predictions, match, so perturbative quantum correlators agree when $B=F$ is enforced. The unified packaging; gravity $R$ and gauge $F$ are components of one curvature $\mathcal{F}$ of one connection $\mathcal{A}$ on one bundle $P_{\mathrm{tot}}$.
Theorem 2
Let $(M,h)$ be a real Lorentzian 44–manifold admitting a complexification $M_{\mathbb{C}}$. Let $P_{\mathrm{tot}}\to M$ be a principal HH–bundle with $H=\mathrm{Spin}(1,3)\times G$ where $G$ is compact, reductive, and contains the SM gauge group. Consider fields $\big(h_{\mu\nu},\mathcal{A}=(\omega,A_{\mathrm{YM}}),\Psi,\Phi,g_{\mu\nu}\big)$ with:
$g_{\mu\nu}=h_{\mu\nu}+i\,B_{\mu\nu}\quad\text{(Hermitian packaging on $M_{\mathbb{C}}$)}.$ (66)
Assume the dynamics are invariant under the automorphism group:
$\mathrm{Aut}(P_{\mathrm{tot}})=\mathrm{Diff}(M)\ltimes\mathrm{Gau}(P_{\mathrm{Spin}}\times_{M}P_{G}),$ (67)
implemented off–shell by a nilpotent BRST differential $s$ (Noether II setting). The Lagrangian density is a diffeo and gauge–natural polynomial in the fields and their covariant derivatives, and may depend on the covariant d’Alembertian $D^{2}$ only through an entire functional calculus $F(D^{2}/M_{*}^{2})$ that commutes with $s$ and all covariant derivatives. In the limit $M_{*}\to\infty$, the quadratic operators for $(h,A,\Psi,\Phi)$ reduce to second order in derivatives, with indices raised and lowered solely by $h$. Reflection positivity on the Euclidean slice as no extra propagating poles other than those of $(h,A,\Psi,\Phi)$. No additional massless higher–spin gauge fields beyond spin–2 $h$ and spin–1 connections; interactions marginal or relevant in the IR. The local BRST anomaly cohomology $H^{1,4}(s\mid d)$ vanishes for the chosen matter representation, no gauge or diffeo anomalies.
Then, modulo ss–exact terms, total derivatives, and higher–dimension operators suppressed by $M_{*}$, the most general action is, after field redefinitions, equivalent to:
[…]
with the following consequences; BRST covariance implies the antisymmetric piece $B_{\mu\nu}$ transforms as a curvature two–form, hence on the real slice:
$B_{\mu\nu}\;\equiv\;(F_{\mathrm{YM}})_{\mu\nu}\quad\text{(in an appropriate adjoint/trace)}\,$, (69)
so $g_{\mu\nu}=h_{\mu\nu}+iB_{\mu\nu}$ is a kinematic packaging and does not introduce an independent 2-form gauge sector. Preserving the full local symmetry off–shell to all loops restricts UV softening to entire functions $F(D^{2}/M_{*}^{2})$ at covariant kinetic operators and consistently at vertices. This renders perturbation theory UV–finite while the IR ($M_{*}\to\infty$) recovers local EYM++matter. Noether II yields the unified Bianchi and Slavnov–Taylor identities, and:
$D_{\mu}J^{\mu}_{a}=0,\qquad\nabla_{\mu}T^{\mu\nu}=F^{\nu\rho}_{a}\,J^{a}_{\rho},$ (70)
with $T^{\mu\nu}$ the Belinfante tensor built from $h$. Up to $S_{\mathrm{top}}$, Euler, Pontryagin, $\theta$–terms, no further diffeo or gauge–natural, unitary, second–order couplings exist.
In this section, we specify the geometric data $P_{\mathrm{tot}}=P_{\mathrm{Spin}}\times_{M}P_{G}$ with structure group $H=\mathrm{Spin}(1,3)\times G$, a single HH–connection $\mathcal{A}=(\omega,A)$ with unified curvature $\mathcal{F}=(R,F)$, and a Lorentzian metric $h$. The master action Eq. (7) is written and shown to be $\mathrm{Diff}(M)\times H$–invariant by construction. From the single bundle identity $D_{\mathcal{A}}\mathcal{F}=0$ we then derive the split into the Riemann and Yang–Mills Bianchi identities, $D_{\omega}R=0$ and $D_{A}F=0$, thereby establishing the unified Noether–II content. Variation with respect to $A$ yields the Yang–Mills equations $D_{A}(*_{h}F)=J_{\rm matter}$ Eq. (10), while variation with respect to $h$ yields Einstein’s equations sourced by the Yang–Mills and matter stress tensors Eq. (9). Therefore a single invariant action produces both sectors’ dynamics and their identities, exactly in Noether’s sense.
Within these principles there are no additional diffeomorphism- and gauge-natural, unitary second-order couplings in four dimensions beyond Einstein–Hilbert, Yang–Mills, and minimal matter terms as anything else is either purely topological, Euler, Pontryagin, $\theta$–terms or higher-dimension and therefore suppressed. Attempts to add an independent antisymmetric metric mode, extra massless higher-spin fields, or non-natural derivative couplings either break the symmetry structure, introduce ghosts or extra poles, or run afoul of anomaly constraints. Thus, in this construction, SM ++ GR is not an arbitrary choice but the unique possibility consistent with the Einstein–Noether program, and the Hermitian geometry makes that uniqueness geometrically explicit via $g_{[\mu\nu]}=F_{\mu\nu}$.
5 Field Geometry and Global Spatial Topology
In HUFT, all fields live on the product principal bundle:
$P_{\mathrm{tot}}\;=\;P_{\mathrm{Spin}}\times_{M}P_{G},\qquad\text{with structure group},\qquad H=\mathrm{Spin}(1,3)\times G,$ (71)
with a single HH–connection $\mathcal{A}=(\omega,A)$ whose curvature $\mathcal{F}=d\mathcal{A}+\mathcal{A}\!\wedge\!\mathcal{A}=(R,F)$ splits into spacetime curvature $R$ and Yang–Mills curvature $F$. On the complexified tangent bundle, we package the metric as a Hermitian field $g=h+iB$. On the real slice $y=0$ used for physics, the antisymmetric piece $B$ is not independent as it is identified with the gauge curvature two-form, $B\equiv F$, so the gauge-field shape is literally the two-form curvature living in the internal fiber.
The internal fiber and vacuum manifold comes from us adopting the minimal simple gauge group:
$G=\mathrm{SU}(5),\qquad\dim G=24,$ (72)
whose maximal torus or Cartan is $T^{4}=\mathrm{U}(1)^{4}$. The electric and hypercharge assignments live along these four commuting directions, the weights of the Cartan.
Spontaneous symmetry breaking proceeds in two steps:
(GUT) $\mathrm{SU}(5)\;\longrightarrow\;\mathrm{SU}(3)_{c}\times\mathrm{SU}(2)_{L}\times\mathrm{U}(1)_{Y},$
(EW) $\mathrm{SU}(2)_{L}\times\mathrm{U}(1)_{Y}\;\longrightarrow\;\mathrm{U}(1)_{\mathrm{EM}}.$
The GUT vacuum manifold is the flag space:
$\mathcal{M}_{\mathrm{GUT}}\;=\;\frac{\mathrm{SU}(5)}{\mathrm{SU}(3)\times\mathrm{SU}(2)\times\mathrm{U}(1)},\qquad\dim\mathcal{M}_{\mathrm{GUT}}=12,$ (73)
and the electroweak vacuum manifold is:
$\mathcal{M}_{\mathrm{EW}}\;\simeq\;\frac{\mathrm{SU}(2)\times\mathrm{U}(1)}{\mathrm{U}(1)_{\mathrm{EM}}}\;\cong\;S^{3}.$ (74)
These spaces encode which gauge directions are unbroken versus broken, massless vs massive respectively.
Throughout this paper, we take spatial slices to be:
$\Sigma_{3}\;\cong\;\mathbb{R}^{3},$ (75)
such as spatially flat and simply connected, in line with current cosmological evidence for near-flat geometry. This choice fixes boundary conditions and mode expansions but does not affect the local variational derivation of the field equations or the unification mechanism.
Other constant-curvature topologies remain compatible with the framework such as a 3-torus $T^{3}$, periodic identifications on three cycles lead to discrete comoving momenta and possible global Aharonov–Bohm phases for gauge holonomies around non-contractible loops. Defect classification uses the same homotopy groups, but spatial infinity is replaced by large embedded 2-tori, 2-spheres, or by cycle representatives. A 3-sphere $S^{3}$, finite spatial volume, harmonic analysis uses $S^{3}$ eigenmodes. Large-radius $S^{2}\subset S^{3}$ still serves to define monopole charges; instanton number is unchanged since it lives in $\pi_{3}(G)$. But in all cases, the internal geometry, fiber $G$, Cartan $T^{4}$, vacuum manifolds, and the identification $B\equiv F$ on the real slice are unchanged, as only global boundary conditions and spectral discreteness differ.
Gauge- and BRST-covariant entire-function regulators $F(D^{2}/M_{*}^{2})$ act as an ultraviolet texture filter, exponentially damping modes with $|p|\gg M_{*}$ without introducing new poles. Operationally, correlators are band-limited on scales shorter than $M_{*}^{-1}$, so fine-grained field wrinkles are smoothed while locality, gauge identities, and unitarity are preserved.
So in HUFT the universe’s field shape in this framework is a single connection on $P_{\mathrm{Spin}}\times_{M}P_{G}$ whose curvature splits into gravity and gauge parts, an internal $\mathrm{SU}(5)$ fiber containing a Cartan $T^{4}$ that organizes charges, vacuum manifolds $\mathcal{M}_{\mathrm{GUT}}$; flag space and $\mathcal{M}_{\mathrm{EW}}\cong S^{3}$ that determine massless vs. massive directions, topological sectors classified by maps from $S^{2}$ and $S^{3}$ into these spaces, and UV-smooth field textures controlled by $M_{*}$. For definiteness, we fix $\Sigma_{3}\simeq\mathbb{R}^{3}$, while noting that $T^{3}$ or $S^{3}$ spatial topologies are also compatible and leave the unification structure intact.
In the present formulation, HUFT requires that the real slice $M$ be an oriented, time-orientable spin 4–manifold supporting the product principal bundle $P_{\mathrm{tot}}=P_{\mathrm{Spin}}\times_{M}P_{G}$ with a single $H=\mathrm{Spin}(1,3)\times G$ connection $\mathcal{A}=(\omega,A)$ and curvature $\mathcal{F}=R\oplus F$. On the real slice, the antisymmetric metric piece satisfies $B\equiv F$, so the gauge sector is encoded directly in the fiber curvature. Chern–Weil matching and anomaly cancellation fix the normalization of $\langle\!\langle F\wedge F\rangle\!\rangle$ and select the minimal internal group $G=\mathrm{SU}(5)$. These are global bundle constraints on $M$, but they do not determine the global topology of spacelike slices $\Sigma_{t}$.
Accordingly, we do not commit the theory to a unique $\Sigma_{t}$ but common choices such as $\mathbb{R}^{3}$, $T^{3}$, or $S^{3}$ are all compatible, provided they admit the required spin structure and principal $G$-bundle. Physical consequences of different $\Sigma_{t}$ arise via boundary conditions such as discrete momenta on $T^{3}$, finite-volume harmonics on $S^{3}$, and possible global holonomies, not from the unification mechanism itself. Any further selection among allowed topologies would have to come from cosmological initial data or an extension of the action that weights topology classes.
6 Quantum Mechanics and Quantum Field Theory from HUFT
Now, to show compatibility with quantum mechanics, we will show the Born rule, Veltman condition, and Dirac equation come out naturally from fibre bundles in HUFT [36, 37, 38]. First, recall that $(M,g_{\mu\nu})$ is the complex Hermitian spacetime of HUFT with real slice $M\subset M_{\mathbb{C}}$. Matter fields live in a complex Hermitian vector bundle:
$\pi:E\to M,\qquad\text{rank}(E)=d\in\mathbb{N},$ (76)
associated with a principal bundle with structure group $G\times U(1)$. On the real slice, there is a positive Hermitian fiber metric:
$h_{x}:E_{x}\times E_{x}\to\mathbb{C},\quad h_{x}(\cdot,\cdot)=\langle\cdot,\cdot\rangle_{x},$ (77)
and a unitary connection $D=\mathrm{d}+{\cal A}$ compatible with $h$. A pure quantum state at $x\in M$ is a ray $\psi_{x}]\in\mathbb{P}(E_{x})$, such as a nonzero vector modulo the $U(1)$ fiber phase. A ray $\psi]\in\mathbb{P}(E_{x})$ is the $1$-dimensional subspace $\{\lambda\psi:\lambda\in\mathbb{C}^{\times}\}$; physical predictions depend only on $\psi]$, equivalently on $\Pi_{[\psi]}=|\psi\rangle\langle\psi|/\langle\psi,\psi\rangle$. Observable $A$ is a Hermitian bundle endomorphism $A\in\Gamma(\mathrm{End}(E))$ at $x$ it has the spectral resolution $A_{x}=\sum_{i}a_{i}\,\Pi_{i,x}$ with orthogonal projectors $\Pi_{i,x}\in\mathrm{End}(E_{x})$.
Our goal is to construct a probability assignment with $\mu_{x}:\{\text{projectors on }E_{x}\}\to[0,1]$ meaning assign a probability number to every projector or outcome at $x$, such that for the outcome subspace $\mathrm{Im}\,\Pi_{i,x}$ we have:
$\mu_{x}(\Pi_{i,x}\,|\,[\psi_{x}])\;=\;\|\Pi_{i,x}\psi_{x}\|_{h_{x}}^{2}/\|\psi_{x}\|_{h_{x}}^{2}.$ (78)
[…]
For $\dim E_{x}\geq 3$, by Gleason’s theorem, the existence of a positive trace-one operator $\rho_{[\psi]}$ on $E_{x}$ such that for every projector $P\in\mathcal{P}$:
$\mu(P\,|\,\psi)=\mathrm{Tr}(\rho_{[\psi]}P).$ (79)
ray invariance and $U(1)$-equivariance of $E$ force $\rho_{[\psi]}$ to be a rank-one projector: $\rho_{[\psi]}=\frac{|\psi\rangle\langle\psi|}{\langle\psi,\psi\rangle}.$ This shows:
$\mu(P\,|\,\psi)=\frac{\langle\psi,P\psi\rangle}{\langle\psi,\psi\rangle}.$ (80)
Taking $P=\Pi_{i}$ gives the Born rule:
$\mu(\Pi_{i}\,|\,\psi)=\|\Pi_{i}\psi\|^{2}/\|\psi\|^{2}.$ (81)
When $\dim E_{x}=2$, we replace A3 by $\sigma$-additivity for measurable fields of positive effects summing to $\mathbf{1}$ positive operator–valued measures (POVMs). By the Busch–Gleason extension [37], we again obtain: $\mu(E\,|\,\psi)=\mathrm{Tr}(\rho_{[\psi]}E)$ for all effects $E$, hence the same Born rule for projectors.
The preceding is fiberwise. To obtain spacetime probabilities for configuration-space localized outcomes, we must fix a Cauchy hypersurface $\Sigma\subset M$ with induced positive measure $\mathrm{d}\Sigma\sqrt{\gamma}$ from the real-slice metric. For a Hermitian line subbundle $L\subset E$ such as a position or detector mode like a localized wave-packet subbundle defined by a smooth section basis, the probability density is the fiber norm induced by $h$:
$\rho(x)=\frac{h_{x}(\Pi_{L}\psi_{x},\Pi_{L}\psi_{x})}{\int_{\Sigma}h_{y}(\psi_{y},\psi_{y})\,\mathrm{d}\Sigma_{y}\sqrt{\gamma(y)}}.$ (82)
Gauge covariance under the $U(1)$ factor of the HUFT structure group is guaranteed because $h$ and $D$ are unitary, $h(\psi,\psi)$ is $U(1)$-invariant, and parallel transport preserves the norm. Here $h(\psi,\psi)\equiv h_{x}(\psi_{x},\psi_{x})$ denotes the Hermitian fiber norm squared of $\psi$ at $x$, in a local frame $(h_{i\bar{j}})$ with $\psi=\psi^{i}e_{i}$ we have $h(\psi,\psi)=\bar{\psi}^{\bar{j}}h_{i\bar{j}}\psi^{i}$. Thus, the only $U(1)$-invariant quadratic functional compatible is the squared $h$-norm.
We let $E_{x}=E^{S}_{x}\otimes E^{E}_{x}$ describe system and environment fibers. For a Schmidt state $|\psi\rangle=\sum_{k}\sqrt{p_{k}}\,|k\rangle_{S}\otimes|k\rangle_{E}$, HUFT’s unitary fiber symmetries contain phase twirls $U_{S}\otimes U_{E}^{\dagger}$ that leave $|\psi\rangle$ invariant called environment-assisted invariance [38]. Equal-weight cases, $p_{k}$ equal force equiprobability on the system outcomes by symmetry, rational weights follow by refinement, and continuity from (A3) yields $\mu(\Pi_{k}\,|\,\psi)=p_{k}=\|\Pi_{k}\psi\|^{2}/\|\psi\|^{2}.$ Within HUFT, states are rays in a Hermitian vector bundle, and measurements are Hermitian bundle endomorphisms. Assuming ray invariance, additivity on orthogonal outcomes, and locality or noncontextuality for subspaces, the unique probability assignment is:
$\Pr(a_{i}|\psi)=\frac{\langle\psi,\Pi_{i}\psi\rangle}{\langle\psi,\psi\rangle}=\frac{\|\Pi_{i}\psi\|_{h}^{2}}{\|\psi\|_{h}^{2}},$ (83)
fiberwise at each spacetime point, and its spacetime version is obtained by integrating the $h$-norm density over the appropriate hypersurface measure from the real-slice metric. This is the Born rule in HUFT’s fiber-bundle language.
The geometric set-up as before holds, but we as well assume a ${\rm Spin}^{c}$ structure to derive the Dirac equation. We let $S\to M$ be the complex spinor bundle with Hermitian fiber metric $h_{S}$ and Clifford map $c:T^{*}M\to{\rm End}(S)$, $c(\alpha)c(\beta)+c(\beta)c(\alpha)=2\,h^{\mu\nu}\alpha_{\mu}\beta_{\nu}\mathbf{1}$, meaning $c$ is a map $c:T^{\ast}M\to\mathrm{End}(S)$. For any one-forms $\alpha,\beta$:
$c(\alpha)c(\beta)+c(\beta)c(\alpha)=2\,h^{\mu\nu}\alpha_{\mu}\beta_{\nu}\,\mathbf{1}.$ (84)
Equivalently, with $c(dx^{\mu})=\gamma^{\mu}$, this reads $\{\gamma^{\mu},\gamma^{\nu}\}=2\,h^{\mu\nu}\mathbf{1}$, as this is the gamma-matrix anticommutation relation. For the connection we let $\nabla^{\rm LC}$ be the Levi–Civita connection of $h_{\mu\nu}$, $\omega^{ab}{}_{\mu}$ its spin connection, and $A_{\mu}\equiv A_{\mu}^{A}T_{A}+A^{(1)}_{\mu}$ the $G\times U(1)$ gauge potential. The total unitary covariant derivative on sections of $S\otimes E_{\rm rep}$ is:
$D_{\mu}=\partial_{\mu}+\tfrac{1}{4}\,\omega^{ab}{}_{\mu}\gamma_{ab}+i\,A^{(1)}_{\mu}\,\mathbf{1}+i\,A_{\mu}^{A}T_{A},\quad\text{with}\quad\gamma_{\mu}\equiv c(e_{\mu}).$ (85)
Compatibility means $D_{\mu}h_{S}=0$ and $D_{\mu},c(e_{\nu})]=\Gamma^{\rho}_{\mu\nu}c(e_{\rho})$, this is equivalent to metric or Clifford compatibility and ensures $D_{\mu}h_{S}=0$.. We consider the Dirac Lagrangian density on the real slice [3, 29]:
[…]
for $\psi\in\Gamma(S\otimes E_{\rm rep})$, this read as
$\psi$ is a section of the tensor–product bundle $S\otimes E_{\rm rep}$, meaning at each spacetime point at $x\in M$:
$\psi(x)\in S_{x}\otimes(E_{\text{rep}})_{x},$ (87)
so $\psi$ is a spinor field from $S$ that as well carries an internal gauge index from $E_{\rm rep}$. Here $\Gamma(\cdot)$ denotes the space of smooth sections. In components, one can think of $\psi^{ai}(x)$ with spinor index $a$ and internal representation index $i$. Varying with respect to $\bar{\psi}$ and using metric or Clifford compatibility gives:
[…]
such as:
$i\,\gamma^{\mu}\Big(\partial_{\mu}+\tfrac{1}{4}\omega^{ab}{}_{\mu}\gamma_{ab}+i\,A^{(1)}_{\mu}+i\,A_{\mu}^{A}T_{A}\Big)\psi\;-\;m\,\psi\;=\;0.$ (89)
This means the Dirac equation is precisely the covariant, unitary, Clifford-compatible section equation on the spinor bundle determined by HUFT’s geometry and gauge structure. But on the complex slice, we would use the holomorphic or anti-holomorphic split, but restriction to $M$ yields the above.
Now we consider the one-loop effective action around slowly varying backgrounds. Let $\Phi$ denote the real components of the Higgs doublet $H$, gauge and ghost fields, and fermions. The quadratic fluctuation operator has a Laplace type on each associated bundle:
$\Delta\;=\;-\,D^{2}\;+\;\mathcal{E},\qquad D^{2}\equiv h^{\mu\nu}D_{\mu}D_{\nu},$ (90)
where $\mathcal{E}\in\Gamma({\rm End}(\mathcal{V}))$ is an endomorphism built from background fields such as Higgs potential curvature, Yukawa endomorphisms, covariant curvatures of $D_{\mu}$, and $\mathcal{V}$ is the vector bundle that the fluctuation field lives in. The regulated one-loop effective action can be written via the heat kernel as [13, 14]:
$\Gamma^{(1)}\sim\tfrac{1}{2}\,{\rm STr}\,\log\Delta\;\propto\;\int_{M}\sqrt{|h|}\;\Big[\frac{\Lambda^{4}}{(4\pi)^{2}}\,a_{0}\;+\;\frac{\Lambda^{2}}{(4\pi)^{2}}\,a_{1}\;+\;\cdots\Big],$ (91)
with Seeley–DeWitt coefficients $a_{k}$ that are local, gauge- and diffeo-invariant fiber traces. The quadratic divergence is controlled by:
$a_{1}\;=\;{\rm STr}\,\mathcal{E}$, (92)
a supertrace over all fluctuating species, bosons with $(+)$ sign, fermions with ()(-) sign, including ghosts.
Specializing to backgrounds where only $H$ is nonzero and slowly varying, $\mathcal{E}$ restricted to each species reduces at leading order to a mass-squared endomorphism that is affine in $H^{\dagger}H$:
$\mathcal{E}\big|_{\text{species}}\;=\;m_{0}^{2}\mathbf{1}\;+\;c_{\text{sp}}\,\lambda_{\text{sp}}\,(H^{\dagger}H)\,\mathbf{1}\;+\;\cdots,$ (93)
with $c_{\text{sp}}$ a group-theory or counting factor and $\lambda_{\text{sp}}$ the relevant coupling. The coefficient of the induced operator $H^{\dagger}H$ in the quadratically divergent part of $\Gamma^{(1)}$ is proportional to the fiberwise supertrace:
${\cal C}_{H^{\dagger}H}\;\;=\;\;{\rm STr}\,\big[\partial_{H^{\dagger}H}\mathcal{E}\big]\;\;=\;\;6\,\lambda\;+\;\frac{9}{4}\,g^{2}\;+\;\frac{3}{4}\,g^{\prime 2}\;-\;6\,y_{t}^{2}\;+\;\cdots,$ (94)
where the displayed terms are, respectively, the Higgs self-coupling, ${\rm SU}(2)$ and ${\rm U}(1)_{Y}$ gauge couplings, and the top Yukawa, the ellipsis denotes smaller Yukawas and any additional fields in the chosen fiber content.
The Veltman condition is treated as the fiber-geometric statement that this supertrace vanishes at some renormalization scale $\mu$:
${\rm STr}\,\big[\partial_{H^{\dagger}H}\mathcal{E}\big](\mu)\;=\;0\quad\Longleftrightarrow\quad 6\,\lambda(\mu)+\frac{9}{4}g^{2}(\mu)+\frac{3}{4}g^{\prime 2}(\mu)-6\,y_{t}^{2}(\mu)+\cdots=0.$ (95)
The scale $\mu$ where ${\rm STr}\,\partial_{H^{\dagger}H}\mathcal{E}(\mu)=0$ is scheme-dependent so the supertrace structure itself is the universal $a_{1}$ coefficient. The interpretation of this is $a_{1}\propto{\rm STr}\,\mathcal{E}$ is a bundle trace of an endomorphism, so the sum rule is just the statement that the quadratic counterterm to $H^{\dagger}H$ disappears when the supertrace over all associated bundles, with correct statistics and ghost structure, is zero. In HUFT with entire-function regulators, quadratic divergences are tamed, but the same $a_{1}$ coefficient governs the finite threshold correction replacing the naive $\Lambda^{2}$ piece, so the supertrace relation remains the geometric criterion for suppressing the Higgs mass renormalization.
From the same geometric structure as before, we define a single holomorphic principal bundle [44]:
$\pi:\ \mathcal{Q}\to M,\qquad\mathrm{Struct}(\mathcal{Q})=\mathcal{H}:=\mathrm{SL}(2,\mathbb{C})_{\mathrm{Lor}}\times G,$ (96)
equipped with a holomorphic connection $\mathcal{A}$ and curvature $\mathcal{F}:=\mathrm{d}\mathcal{A}+\mathcal{A}\wedge\mathcal{A}$. We use the Cartan block decomposition:
$\mathcal{A}=\omega\oplus A,\qquad\mathcal{F}=R\oplus F,$ (97)
where $\omega$ is the complex spin connection, $R$ the complexified Riemann curvature, and $A,F$ the internal gauge connection and curvature. We let $e$ denote the complex soldering form; on the real slice $e$ induces $h=g_{(\mu\nu)}$.
We now choose an $\mathrm{Ad}$-invariant, nondegenerate bilinear form on the Lie algebra:
$\langle X,Y\rangle_{\mathcal{H}}:=\kappa_{\mathrm{Lor}}\operatorname{Tr}_{\mathrm{Lor}}(X_{\mathrm{Lor}}Y_{\mathrm{Lor}})+\kappa_{G}\,\operatorname{Tr}_{G}(X_{G}Y_{G}),\qquad X=X_{\mathrm{Lor}}\oplus X_{G},$ (98)
with $\operatorname{Tr}$ the holomorphically normalized Killing forms on each factor. We fix the ratio $\rho:=\kappa_{\mathrm{Lor}}/\kappa_{G}$ by the Chern–Weil matching condition:
$\int_{M}\operatorname{Tr}_{\mathcal{H}}(\mathcal{F}\wedge\mathcal{F})\ =\ \int_{M}\big(\kappa_{\mathrm{Lor}}\operatorname{Tr}_{\mathrm{Lor}}(R\wedge R)+\kappa_{G}\operatorname{Tr}_{G}(F\wedge F)\big)\ \in\ 8\pi^{2}\,\mathbb{Z}$, (99)
compatible with anomaly cancellation and boundary conditions specified below. With this normalization, we define the one-coupling holomorphic action:
$S_{\mathrm{HUFT}}=\Re\int_{M}\Bigg[\frac{1}{g_{\star}^{2}}\,\langle\mathcal{F},\star\mathcal{F}\rangle_{\mathcal{H}}\ +\ \lambda\,\langle B,\mathcal{F}\rangle_{\mathcal{H}}\ +\ \mu\,\langle e\wedge e,\ R\rangle_{\mathrm{Lor}}\ +\ S_{\mathrm{matter}}(\mathcal{A},e;\,\Psi)\Bigg],$ (100)
where $B$ is an auxiliary holomorphic $\mathfrak{h}$-valued two-form and $\star$ is the complex Hodge operator. Varying (100) with respect to $B$ gives:
$\delta_{B}S:\qquad B=\frac{1}{\lambda}\,\star\mathcal{F}\quad\Rightarrow\quad B=\mathcal{F}\ \ \text{on the real slice},$ (101)
where the last statement follows after imposing reality or Hodge conditions on $(M,h)$. Eliminating $B$ yields:
$S_{\mathrm{HUFT}}=\Re\int_{M}\Big[\frac{1}{g_{\star}^{2}}\,\langle\mathcal{F},\star\mathcal{F}\rangle_{\mathcal{H}}+\mu\,\langle e\wedge e,\ R\rangle_{\mathrm{Lor}}+S_{\mathrm{matter}}\Big].$ (102)
Varying with respect to $\omega$ and $e$ gives, on the real slice, metric compatibility and torsionlessness:
$\nabla^{\mathrm{LC}}_{\mu}h_{\alpha\beta}=0,\qquad T(e,\omega)=0,$ (103)
and the Einstein equation sourced by the gauge–matter stress tensor. Variation with respect to $A$ gives the Yang–Mills equations with respect to $h$. Writing (102) on $(M,h)$:
$S_{\mathrm{slice}}=\int_{M}\Big[\frac{1}{16\pi G_{N}}\,R(h)\,\mathrm{vol}_{h}+\frac{1}{4g^{2}}\,\operatorname{Tr}_{G}(F_{\mu\nu}F^{\mu\nu})\,\mathrm{vol}_{h}+\mathcal{L}_{\mathrm{matter}}\Big],$ (104)
with:
$\frac{1}{16\pi G_{N}}:\frac{1}{4g^{2}}\ =\ \mu\,\kappa_{\mathrm{Lor}}:\frac{1}{g_{\star}^{2}}\,\kappa_{G}\ \ \Longrightarrow\ \ \frac{G_{N}}{g^{2}}\ =\ \frac{g_{\star}^{2}}{4\pi}\,\frac{\kappa_{G}}{\mu\,\kappa_{\mathrm{Lor}}}\ =\ \frac{g_{\star}^{2}}{4\pi}\,\frac{1}{\mu\,\rho}.$ (105)
The relative gravity–gauge normalization is not a free dial as it is set by the geometric ratio $\rho=\kappa_{\mathrm{Lor}}/\kappa_{G}$ fixed in (99) and by the universal constants $(g_{\star},\mu)$ entering (100).
The entire-function regulator $f(\Box/M_{*}^{2})$ preserves the holomorphic Ward and Slavnov–Taylor identities [8, 9, 10], yielding a common renormalization above $M_{*}$:
$\beta_{\mathcal{H}}(\mu>M_{*})=0\qquad\Rightarrow\qquad g_{\mathrm{Lor}}(\mu)=g_{G}(\mu)=g_{\star}.$ (106)
Combined with (105), this gives a UV plateau with locked gravity–gauge normalization inherited by threshold matching to the IR.
To show the internal gauge symmetries from our fibre bundle we let $E\to M$ be a rank-$n$ holomorphic Hermitian vector bundle for matter, with $c_{1}(E)=0$, a nowhere-vanishing holomorphic volume form $\Omega_{E}\in H^{0}(M,\wedge^{n}E^{*})$, and assume $E$ is slope-stable so that the compatible connection is Hermitian–Yang–Mills. Preservation of $(h_{E},\Omega_{E})$ reduces the internal structure group to $\mathrm{SU}(n)$, and the internal connection has holonomy contained in $\mathrm{SU}(n)$.
We demand a single adjoint reduction to the real-slice group, correct hypercharge quantization obtained from an integral pairing derived from (99), and chiral, anomaly-free matter for one family from associated bundles of $E$. Then the minimal rank is $n=5$.
Now for the mixed Ward identities to show the Diff $\leftrightarrow$ Gauge interlock we fix holomorphic gauge conditions for $\mathcal{A}$ and diffeomorphisms on $M$. To preserve the gauge conditions, an infinitesimal diffeomorphism $\delta_{\xi}$ must be accompanied by a compensating internal transformation $\delta_{\epsilon}$:
$\delta_{\xi}\mathcal{A}+\delta_{\epsilon}\mathcal{A}=0\quad\text{in gauge},\qquad\epsilon=\epsilon[\xi;\,\mathcal{A},e].$ (107)
The associated Ward identity for connected correlators $\langle\cdots\rangle$ is:
$\partial_{\mu}\langle T^{\mu\nu}(x)\,\mathcal{O}(y)\rangle=\alpha(\rho)\,\partial_{\mu}\langle J^{\mu}_{\mathrm{gauge}}(x)\,\mathcal{O}(y)\rangle+\cdots,\qquad\alpha(\rho)\propto\frac{\kappa_{G}}{\kappa_{\mathrm{Lor}}}=\rho^{-1},$ (108)
exhibiting an explicit cross-sector conservation relation with coefficient fixed by the geometric ratio $\rho$.
Finally, we will derive the Ehrenfest theorem from the unified geometry on the real slice $(M,h)$. We fix a global time function $t$ with Cauchy slices $\Sigma_{t}$, unit normal $n^{\mu}$, induced metric $\gamma_{ij}$, and measure $d\Sigma\,\sqrt{\gamma}$. Let matter states be sections $\psi$ of the Hermitian bundle with hh–inner product and minimal coupling through the unitary covariant derivative:
$D_{\mu}\;=\;\partial_{\mu}+\tfrac{1}{4}\omega^{ab}{}_{\mu}\gamma_{ab}+i\,A_{\mu},\qquad[D_{\mu},D_{\nu}]\;=\;i\,F_{\mu\nu}.$ (109)
We define the kinetic momentum operator $\hat{\Pi}_{\mu}:=-i\hbar D_{\mu}$, the position operator $\hat{X}^{\mu}$, multiplication by the coordinate function on $\Sigma_{t}$, and for any bundle endomorphism observable $\hat{O}$ define the expectation value on $\Sigma_{t}$ by:
$\langle\hat{O}\rangle_{t}\;:=\;\frac{\displaystyle\int_{\Sigma_{t}}\!\sqrt{\gamma}\;\psi^{\dagger}\hat{O}\,\psi}{\displaystyle\int_{\Sigma_{t}}\!\sqrt{\gamma}\;\psi^{\dagger}\psi}.$ (110)
We take dynamics to be generated by the minimally coupled covariant Hamiltonian $\hat{H}$ either Dirac, Pauli, or Schrödinger built from $h$ and $D_{\mu}$, so that $i\hbar\,\partial_{t}\psi=\hat{H}\psi$ and the probability current is conserved.
We assume standard falloff, so boundary terms at spatial infinity vanish, and take $V$ a real scalar section such as a gauge singlet potential. Then the expectation values of position and kinetic momentum obey:
$\frac{D}{dt}\,\big\langle\hat{X}^{\mu}\big\rangle$ $=\frac{1}{m}\,\big\langle\hat{\Pi}^{\mu}\big\rangle,$ (111)
$\frac{D}{dt}\,\big\langle\hat{\Pi}_{\mu}\big\rangle$ $=-\,\big\langle\nabla_{\mu}V\big\rangle+\big\langle q\,F_{\mu\nu}\,\dot{X}^{\nu}\big\rangle-\big\langle\Gamma^{\rho}_{\mu\nu}\,\hat{\Pi}_{\rho}\,\dot{X}^{\nu}\big\rangle,$ (112)
where $\Gamma^{\rho}_{\mu\nu}$ is the Levi–Civita connection of $h$, the overdot denotes contraction with $n^{\mu}$ as the foliation flow, and $F_{\mu\nu}$ is the gauge curvature which equals the antisymmetric piece $B_{\mu\nu}$ of the Hermitian metric on the real slice ($B=F$). The derivative $D/dt$ is the Levi–Civita covariant time derivative along the foliation, so that measure and connection effects are included.
Conservation of probability current gives the covariant continuity equation $\partial_{t}(\sqrt{\gamma}\,\psi^{\dagger}\psi)+\nabla_{i}(\sqrt{\gamma}\,j^{i})=0$, so boundary terms from $\partial_{t}\sqrt{\gamma}$ and $\nabla_{i}$ cancel in the time derivative of (110). Thus, we obtain the Heisenberg identity in covariant form:
$\frac{D}{dt}\langle\hat{O}\rangle=\frac{i}{\hbar}\,\langle[\hat{H},\hat{O}]\rangle+\langle\partial_{t}\hat{O}\rangle$, (113)
with the Levi–Civita correction precisely reproducing the $\Gamma$–terms below.
For $\hat{O}=\hat{X}^{\mu}$, minimal coupling implies $\hat{\Pi}_{\alpha},\hat{X}^{\mu}]=-i\hbar\,\delta^{\mu}{}_{\alpha}$, and for the kinetic Hamiltonian $\hat{H}_{\rm kin}=\tfrac{1}{2m}\,h^{\alpha\beta}\hat{\Pi}_{\alpha}\hat{\Pi}_{\beta}$ we obtain:
$\frac{i}{\hbar}[\hat{H}_{\rm kin},\hat{X}^{\mu}]=\frac{1}{2m}\,h^{\alpha\beta}\Big(\hat{\Pi}_{\alpha}[\hat{\Pi}_{\beta},\hat{X}^{\mu}]+[\hat{\Pi}_{\alpha},\hat{X}^{\mu}]\hat{\Pi}_{\beta}\Big)=\frac{1}{m}\,\hat{\Pi}^{\mu},$ (114)
yielding (111) after taking expectation values and accounting for the foliation, no explicit time dependence of $\hat{X}^{\mu}$.
For $\hat{O}=\hat{\Pi}_{\mu}$ we use curvature and metric compatibility and define the kinetic momentum operator:
$\hat{\Pi}_{\mu}=-i\hbar\,D_{\mu},$ (115)
with the kinetic Hamiltonian:
$\hat{H}_{\rm kin}=\frac{1}{2m}h^{\alpha\beta}\hat{\Pi}_{\alpha}\hat{\Pi}_{\beta}.$ (116)
Curvature shows up as a commutator, by definition of the gauge curvature:
$\hat{\Pi}_{\mu},\hat{\Pi}_{\nu}]=-i\hbar\,F_{\mu\nu}.$ (117)
With the metric compatibility:
$\nabla_{\lambda}h_{\mu\nu}=0,$ (118)
meaning $h^{\alpha\beta}$ is covariantly constant, so it commutes with $\hat{\Pi}_{\mu}$ so, no extra terms from derivatives of the metric when you commute things. Then the Heisenberg force operator follows:
$\frac{i}{\hbar}[\hat{H}_{\rm kin},\hat{\Pi}_{\mu}]=\frac{1}{2m}\,h^{\alpha\beta}\Big(\hat{\Pi}_{\alpha}[\hat{\Pi}_{\beta},\hat{\Pi}_{\mu}]+[\hat{\Pi}_{\alpha},\hat{\Pi}_{\mu}]\hat{\Pi}_{\beta}\Big)=\frac{1}{m}\,h^{\alpha\beta}\,\Big(\hat{\Pi}_{\alpha}F_{\beta\mu}+F_{\alpha\mu}\hat{\Pi}_{\beta}\Big),$ (119)
the right-hand side is the symmetrized product of momentum with field strength, so the operator is Hermitian. Whose symmetric part gives the Lorentz-force term $q\,F_{\mu\nu}\,\dot{X}^{\nu}$ in expectation values. The scalar potential contributes $(i/\hbar)[V,\hat{\Pi}_{\mu}]=-\nabla_{\mu}V$. Finally, when transporting the expectation value with the time-dependent volume form and frame, the Levi–Civita piece from $D/dt$ supplies $\langle\Gamma^{\rho}_{\mu\nu}\,\hat{\Pi}_{\rho}\,\dot{X}^{\nu}\rangle$, completing (112). Since on the real slice $B_{\mu\nu}=F_{\mu\nu}$, the gauge-force term is literally the antisymmetric part of the Hermitian metric.
In a local inertial frame at a point $\Gamma=0$ the relations reduce to by taking expectation values and writing:
$\hat{\Pi}^{\nu}:=h^{\nu\alpha}\hat{\Pi}_{\alpha},$ (120)
and
$\hat{v}^{\nu}:=\frac{\hat{\Pi}^{\nu}}{m},$ (121)
gives the Lorentz-force term in Ehrenfest form:
$\frac{d}{dt}\langle\hat{\Pi}_{\mu}\rangle=\left\langle\frac{i}{\hbar}[\hat{H}_{\rm kin},\hat{H}_{\mu}]\right\rangle,\to q\langle F_{\mu\nu}\hat{v}^{\nu}\rangle,$ (122)
in curved space, we add the separate Levi–Civita piece but note that when you take the time derivative of the momentum expectation, you must use the covariant time derivative:
$\frac{D}{dt}\langle\hat{\Pi}_{\mu}\rangle=\frac{d}{dt}\langle\hat{\Pi}_{\mu}\rangle-\langle\Gamma_{\mu\nu}^{\rho}\hat{\Pi}_{\rho}\dot{X}^{\nu}\rangle,$ (123)
When transporting the expectation value, that’s the gravitational part. That extra term is what we call the Levi–Civita piece. It’s the correction needed so that the equation of motion is tensorial coordinate-independent. Without it, you’d be differentiating components as if the basis were fixed, flat space. With it, the quantum expectation values reduce to the geodesic with Lorentz force in the classical limit.
We find the Ehrenfest relations:
$\frac{d}{dt}\langle\hat{X}^{\mu}\rangle=\frac{1}{m}\langle\hat{\Pi}^{\mu}\rangle,\qquad\frac{d}{dt}\langle\hat{\Pi}_{\mu}\rangle=-\langle\nabla_{\mu}V\rangle+\langle q\,F_{\mu\nu}\,\dot{X}^{\nu}\rangle,$ (124)
geodesic drift plus the Lorentz force in expectation.
On a static slice of flat space with coordinates $(t,\mathbf{x})$, $\Gamma=0$ and $h_{\mu\nu}\to\eta_{\mu\nu}$. Writing $\hat{\bm{\pi}}=-i\hbar\,\bm{\nabla}-q\,\mathbf{A}$ and $\hat{\mathbf{v}}=\hat{\bm{\pi}}/m$:
$\frac{d}{dt}\langle\hat{\mathbf{x}}\rangle=\frac{1}{m}\langle\hat{\bm{\pi}}\rangle,\qquad\frac{d}{dt}\langle\hat{\bm{\pi}}\rangle=-\langle\bm{\nabla}V\rangle+q\,\big\langle\mathbf{E}+\hat{\mathbf{v}}\times\mathbf{B}\big\rangle,$ (125)
which collapses to the textbook Ehrenfest theorem when $\mathbf{B}=0$ and $V=V(\mathbf{x})$.
Our observables are bundle endomorphisms compatible with the unitary structure, so this excludes non-covariant ad hoc couplings and ensures the identities above are representation-independent. Gravity enters only through $h$ via $\Gamma$, while gauge forces enter through $F$ but both arise from the same unified connection in the holomorphic theory, so the quantum to classical correspondence is an internal consequence of the geometry, not an additional postulate.
7 Conclusion
We have presented a symmetry–first unification in which gravity and gauge interactions arise from a single geometric framework, a product bundle with one master connection $\mathcal{A}=(\omega,A)$, one curvature $\mathcal{F}=(R,F)$, and a Hermitian field $g=h+iB$ on the complexified spacetime. From a single $\mathrm{Diff}(M)\times G$–invariant action, variation reproduces the Einstein and Yang–Mills equations together with their paired Bianchi identities, realizing the Einstein–Noether program of deriving dynamics and identities from one invariant structure.
Within this unified kinematic and variational structure, there are no additional diffeomorphism- and gauge-natural, unitary, second-order couplings in four dimensions beyond Einstein–Hilbert, Yang–Mills, and minimal matter terms. This makes the SM and GR content a consequence of symmetry and naturalness rather than an arbitrary choice, and the Hermitian packaging renders this uniqueness geometrically and explicitly through $g_{[\mu\nu]}=F_{\mu\nu}$.
Quantum kinematics and measurement descend from the same data. The states are rays in a Hermitian bundle with unitary connection $D$, and the unique probability assignment compatible with our axioms is the Born rule, fiberwise, its spacetime version follows by integrating the hh–norm on Cauchy slices with the induced measure $\sqrt{\gamma}\,d\Sigma$. This ties probabilities to the same real-slice geometry that governs dynamics.
On the perturbative side, implementing entire-function regulators at covariant kinetic operators and vertices yields a UV-finite framework, whose IR limit reproduces local EYM and matter. The unified Noether–II content includes covariant conservation laws and Slavnov–Taylor identities, ensuring consistency of the regulated theory.
Taken together, these results realize a form of geometric unification in which parallel transport, curvature, conserved currents, and measurement all arise from the same $(h,A)$ data, with $B=F$ eliminating superfluous antisymmetric gravitational modes.
Several targeted directions can be follow, a group or metric locking principle either geometric or topological is needed to fix the relative normalization and uniquely determine the internal group, elevating our shared-structure unification toward a strict GUT selection. On the quantum–classical interface, the covariant Ehrenfest lemma, linking the time evolution of expectation values in quantum theory to the classical equations of motion, shows that geodesic drift and Lorentz force emerge in expectation directly from the unified geometry, extending this to open-system dynamics and decoherence on curved backgrounds.
Acknowledgments
We thank Martin Green and Arvin Kouroshnia for helpful discussions. Research at the Perimeter Institute for Theoretical Physics is supported by the Government of Canada through Industry Canada and by the Province of Ontario through the Ministry of Research and Innovation (MRI).